OpenStudy (anonymous):

Determine the soultion of the system. show your work. y=5x- 9 y= 2x+6

OpenStudy (danjs):

solve one for x and substitute that into the second

OpenStudy (danjs):

From equation 1) \[x = \frac{ y+9 }{ 5 }\]

ganeshie8 (ganeshie8):

substituting \(y\) is easy here

OpenStudy (danjs):

so, \[y = 2*[ \frac{ y+9 }{ 5 } ] + 6\]

OpenStudy (danjs):

Solve for Y, and use that value in either equation to find x.

OpenStudy (anonymous):

im lost.. i dont understand any of it

OpenStudy (danjs):

You have 2 equations with 2 unknown variables x and y. You need to find out what values x and y need to be so that both of the equations are true.

OpenStudy (danjs):

So what you can do, as i did above, is solve one equation for x. Then, as i also did above, Substitute that x value you solved for into the next equation. You are left with an equation with one variable. You can solve for that variable.

OpenStudy (danjs):

\[y = 2 * \frac{ y+9 }{ 5 } + 6\]

OpenStudy (danjs):

Multiply everything by 5 to get rid of the fraction. \[5y = 2(y+9) + 30\]

OpenStudy (danjs):

Distribute the 2 into the parenthesis, 5y = 2y + 18 + 30

OpenStudy (danjs):

separate the variable terms and the number constant terms on separate sides, 5y - 2y = 18 + 30

OpenStudy (danjs):

You are left with, 3y = 48

OpenStudy (danjs):

divide both sides of this by 3 y = 48 / 3

OpenStudy (danjs):

So now you have the value for y y = 48 / 3 y = 16

OpenStudy (danjs):

Pick any of the first original equations, y= 5x- 9 y= 2x+6 You already found y = 16 Put y = 16 into either of those equations above to find what the x value is. For example the first equation. 16 = 5x - 9 x = (16 + 9)/5

OpenStudy (triciaal):

another approach y = y set them equal and solve for x substitute this value in either of the original expression for y to get the value y = 5x -9 and y = 2x + 6 5x -9 = 2x + 6 5x - 2x = 6 + 9 3x = 15 x = 5 y = 5(5) - 9 = 25 - 9 = 16 check in eqn 2 16 = 2(5) + 6 = 16 lhs = rhs (5, 16)

OpenStudy (fibonaccichick666):

still lost?